# Iris RESTful Server and IrisTileSource: An Iris implementation for existing OpenSeaDragon viewers

### v1

@@ Abstract
The Iris File Extension (IFE) is a low overhead performance-oriented whole slide image (WSI) file format designed to improve the image rendering experience for pathologists and simplify image management for system administrators. However, static hypertext transfer protocol (HTTP) file servers cannot natively stream subregions of high-resolution image files, such as the IFE. The majority of contemporary WSI viewer systems are designed as browser-based web applications and leverage OpenSeaDragon as the tile-based rendering framework. These systems convert WSI files to Deep Zoom Images (DZI) for compatibility with simple static HTTP file servers. In order to address this limitation, we have developed the Iris RESTful Server, a low-overhead HTTP server with a RESTful API that is natively compatible with the DICOMweb WADO-RS API. Written in C++ with Boost Beast HTTP and Asio networking libraries atop the public IFE libraries, the server offers both security and high performance. Testing shows that a single instance can handle over 5000 tile requests per second with a median latency of 21 ms on a private network. We also developed and merged a new OpenSeaDragon TileSource, compatible with the Iris RESTful API, into the next OpenSeaDragon release, enabling simple and immediate drop-in replacement of DZI images within WSI viewer stacks. Designed as a secure cross-origin resource sharing microservice, this architecture includes detailed deployment instructions for new or existing WSI workflows, and the public examples.restful.irisdigitialpathology.org subdomain is provided as a development tool to accelerate WSI web viewer development.
@@ 1 Introduction
The Iris RESTful Server is written in C++ using the low-level and highly templated Boost Beast HTTP [11] and Asio networking libraries, and is built on the open-source IFE libraries that we have made publicly available[12]. The Iris RESTful Server was designed to be deployed as a modular and scalable containerized microservice accessible via a separate cross-origin resource sharing (CORS) domain to integrate with current IMS WSV applications. We also developed and merged a new OSD TileSource, the IrisTileSource, into the next OSD release so that IFE encoded files can be immediately and seamlessly incorporated into existing digital pathology workflows. Incorporating IrisTileSource requires changing four lines of code within a viewer source (see Code listing 3 below). Finally we make openly available the https://examples.restful.irisdigitalpathology.org/ subdomain, which we refer to as the Iris RESTful Example domain, as a fully Iris RESTful API compliant server with example IFE encoded slides for community use as a tool to aid in streamlining the integration if IFE encoded slides.
@@ 2.1 Iris RESTful Server
The Iris RESTful server was designed as an extremely high-performance, lightweight, and secure dedicated slide-serving microservice. It is configured to scale independently of the IMS and only returns slide tiles in response to Iris RESTful API requests – though the server may be optionally configured to perform static file services such as providing an OSD viewer web-application. The server comprises two independently operating stacks: 1) the networking stack and 2) the file system stack (Figure 2). The server has security logic that parses API requests and discards requests that do not adhere strictly to the WADO-RS or Iris RESTful API (Code listings 1). When the server is configured as a static HTTP web server, it may also stream DZI image tiles for dual IFE and DZI compatibility; however, it will only return known file types and limit the file directory scope to ensure continued system security. Furthermore, the system automatically configures transport layer security (TLS) connections for end-to-end encryption when deployed with application load balancers to ensure Health Insurance Portability and Accountability Act (HIPAA) compliance.
@@ 3.1 Iris RESTful Server
In a more scalable configuration, a cluster service can be deployed using a container orchestrator system such as Kubernetes or Kubernetes-derived services such as those by Amazon Web Service (Elastic Container Service) or Google Cloud (Cloud Run). This design delegates container scaling to a separate hypervisor system, and is more appropriate for large-scale implementations, such as institutions with dedicated hospital information technology offices familiar with scalable container orchestration services. Although static file serving capabilities can be enabled in this configuration (this is the architecture of examples.restful.irisdigitalpathology.org), for this type of production workflow we recommend using the server as a CORS domain where the IMS/PACS service (such as SECTRA AB or Roche) provides the WSV application from their PACS server instance and assigns the Iris-RESTful implementation as the OSD ‘serverUrl’ (see OSD IrisTileSource implementation Section 3.2 below). When deployed in this configuration, the RESTful server automatically generates self-signed 2048-bit RSA certificates that it uses to coordinate a secure connection with the container orchestrator / load balancer to ensure end-to-end encryption, even within a hospital’s secure network or a provider virtual private network, compliant with HIPAA.
@@ 3.2 OSD IrisTileSource
The IrisTileSource can be implemented within existing viewers by substituting only four lines of JS code in the OSD viewer class declaration. Rather than providing an OSD viewer “tileSources” parameter with the path to DZI image file, you must configure the OSD viewer “tileSources” as the new derived IrisTileSource class and assign that class the “serverUrl” parameter with the Iris RESTful server domain as well as the IFE encoded slide file name (excluding the ‘.iris’ extension). See the example HTML in Code listings 3 for simple functional example web-viewer application.
@@
<script>
@@ 4.1 Materials and Methods
Performance was evaluated for both a single-instance local network deployment (Figure 3, right) and the public facing Iris RESTful Example domain in comparison to other well-established and publicly available WSI servers, specifically Path-Presenter and the National Cancer Institute (NCI) imaging data commons (IDC) Slim viewer[6]. Performance was evaluated using Locust[14], a widely accepted open-source and flexible load testing tool. A ramped load test was chosen to allow for automatic instance scaling in the publicly facing WSI servers. Users were scaled up to 160 and 220 concurrent users for public WSI server and local network tests, respectively. Each virtual user issued a tile request at 1-2 ms intervals after receiving the prior response. Custom locust python scripts were created for each target server and tiles were randomly sampled from within image bounds. These scripts have been made available, along with the raw results of these tests (see Data Availability section). All tests were performed on a 2020 13-in M1 MacBook Pro (Apple, Palo Alto, CA) with 8 GB of RAM running macOS 15.4, with Locust scripts executed using Python 3.11 in an Anaconda environment over a $\geq$ 200 mbps wireless (Wifi) connection.
The single-instance local network deployment was hosted on Turing RTK1 (Turing Machines Inc, Cupertino, CA) 8-core Rockchip RK3588 system on module (SoM) circuit board with 8 GB of RAM running Ubuntu 24.04 within a Docker container and with a 1 gigabit Ethernet switch. The Turing RTK1 SoM is similar to the Raspberry Pi CM4 or the Jetson Orin Nx. The publicly facing Iris RESTful Example domain is hosted on AWS using an application load balancer in front of a Fargate elastic container service with 1 virtual CPU and 2 GB of RAM and set to scale-out at either 70% CPU consumption or greater than 500 requests/sec per instance. Both the local network deployment and Iris RESTful Example domain served IFE formatted tiles using the Iris RESTful API. Path-Presenter served DZI formatted tiles using a static HTTP server without an API. The NCI IDC repository served DICOM formatted frames using the WADO-RS API.
@@ 5 Discussion
We previously published on the rapid encoding speeds of IFE files from source WSI files such as DICOM due to the highly parallelized encoding capabilities afforded by the layer-agnostic tile ordering within the file specification[10]. This aids in the value of the Iris RESTful server by allowing rapid conversion from source DICOM to IFE files, which may then be served to OSD viewers. The current versions of the Iris Codec IFE libraries can generate a derived pyramid file (with 2x or 4x downsampling) from a single highest-resolution DICOM or vendor format source within 1-2 mintues – or much faster if OpenSlide decoding[9] is not needed. We continue to work on improving encoding systems and in future releases will add stream encoding (a slide scanner API) to allow for simultaneous generation of archive DICOM and IFE files directly from the scanner to make IFE slides available to Iris RESTful in WSI workflows immediately.
In this paper we performed stressful load testing on WSI systems to assess system durability and performance when these servers. This includes both expected daily loads early in the load test, as well as abnormally high volumes of requests. The early metrics better align with what we have previously observed with the Slim viewer of an approximately 200 ms median response time (Figure 4, left-side) in our prior publication[10]. The Iris RESTful Example domain response times times of approximately 20-30 ms also align closely with what we have experienced when using our example domain during the development of the IrisTileSource and example browser-based WSVs.
@@ 6 Conclusion
The Iris RESTful Server and OSD IrisTileSource allow for immediate and easy replacement of DZI images with IFE encoded slide files within digital pathology WSI workflows. Prepackaged container images allow for server deployment in a scalable microservice architecture and require altering only four lines of viewer code to begin rendering with the new IrisTileSource within existing OSD viewers. Written in C++ with the highly templated Boost Beast library and on top of the IFE deserialization library this is a low-overhead fast deployment that can issue slide tiles in 21 ms on average under an high request load of over 3883 tile requests per second, even when hosted on a small \$200 USD compute module chipboard (the equivalent of a Raspberry-Pi CM4). We provide the Iris RESTful Example domain, a server that hosts example IFE encoded slide files and responds natively to Iris RESTful API calls, for community use to evaluate HTTPS implementations with the Iris RESTful API and accelerate the development process.

### v3

@@ Abstract
The Iris File Extension (IFE) is a low overhead performance-oriented whole slide image (WSI) file format designed to improve the image rendering experience for pathologists and simplify image management for system administrators. However, static hypertext transfer protocol (HTTP) file servers cannot natively stream subregions of high-resolution image files, such as the IFE. The majority of contemporary WSI viewer systems are designed as browser-based web applications and leverage OpenSeaDragon as the tile-based rendering framework. These systems convert WSI files to Deep Zoom Images (DZI) for compatibility with simple static HTTP file servers. To address this limitation, we have developed the Iris RESTful Server, a low-overhead HTTP server with a RESTful API that is natively compatible with the DICOMweb WADO-RS API. Written in C++ with Boost Beast HTTP and Asio networking libraries atop the public IFE libraries, the server offers both security and high performance. Testing shows that a single Raspberry Pi equivalent system can handle a peak of 5,061 req/s (average 3,883 req/s) with a median latency of 21 ms on a private (i.e. hospital) network. We also developed and merged a new OpenSeaDragon TileSource, compatible with the Iris RESTful API, into the next OpenSeaDragon release, enabling simple and immediate drop-in replacement of DZI images within WSI viewer stacks. Designed as a secure cross-origin resource sharing microservice, this architecture includes detailed deployment instructions for new or existing WSI workflows, and the public examples.restful.irisdigitalpathology.org subdomain is provided as a development tool to accelerate WSI web viewer development. All relevant Iris software is available under the open-source MIT software license.
@@ 1 Introduction
The Iris RESTful Server is written in C++ using the low-level and highly templated Boost Beast HTTP [11] and Asio networking libraries, and is built on the open-source IFE libraries that we have made publicly available[12]. The Iris RESTful Server was designed to be deployed as a modular and scalable containerized microservice accessible via a separate cross-origin resource sharing (CORS) domain to integrate with current IMS WSV applications. We also developed and merged a new OSD TileSource, the IrisTileSource, into the next OSD release so that IFE encoded files can be immediately and seamlessly incorporated into existing digital pathology workflows. Incorporating IrisTileSource requires changing four lines of code within a viewer source (see Code listing 3 below). Finally we make openly available the https://examples.restful.irisdigitalpathology.org/ subdomain, which we refer to as the Iris RESTful Example domain, as a fully Iris RESTful API compliant server with example IFE encoded slides for community use as a tool to aid in streamlining the integration of IFE encoded slides.
@@ 2.1 Iris RESTful Server
The Iris RESTful server was designed as an extremely high-performance, lightweight, and secure dedicated slide-serving microservice. It is configured to scale independently of the IMS and only returns slide tiles in response to Iris RESTful API requests – though the server may be optionally configured to perform static file services such as providing an OSD viewer web-application. The server comprises two independently operating stacks: 1) the networking stack and 2) the file system stack (Figure 2). The server has security logic that parses API requests and discards requests that do not adhere strictly to the WADO-RS or Iris RESTful API (Code listings 1). When the server is configured as a static HTTP web server, it may also stream DZI image tiles for dual IFE and DZI compatibility; however, it will only return known file types and limit the file directory scope to ensure continued system security. Furthermore, the system automatically configures transport layer security (TLS) connections for end-to-end encryption when deployed with application load balancers to fulfill the in-transit Health Insurance Portability and Accountability Act (HIPAA) requirement.
@@ 3.1 Iris RESTful Server
In a more scalable configuration, a cloud cluster service can be deployed using a container orchestrator system such as Kubernetes or Kubernetes-derived services such as those by Amazon Web Service (Elastic Container Service) or Google Cloud (Cloud Run). This design delegates container scaling to a separate hypervisor system, and is more appropriate for large-scale implementations, such as institutions with dedicated hospital information technology offices familiar with scalable container orchestration services. Although static file serving capabilities can be enabled in this configuration (this is the architecture of examples.restful.irisdigitalpathology.org), for this type of production workflow we recommend using the server as a CORS domain where the IMS/PACS service (such as SECTRA AB or Roche) provides the WSV application from their PACS server instance and assigns the Iris-RESTful implementation as the OSD ‘serverUrl’ (see OSD IrisTileSource implementation Section 3.2 below). When deployed in this configuration, the RESTful server automatically generates self-signed 2048-bit RSA certificates that it uses to coordinate a secure connection with the container orchestrator / load balancer to ensure end-to-end encryption, even within a hospital’s secure network or a provider virtual private network, fulfilling the in-transient HIPAA encryption requirement.
@@ 3.2 OSD IrisTileSource
The IrisTileSource can be implemented within existing viewers by substituting only four lines of JS code in the OSD viewer class declaration. Rather than providing an OSD viewer “tileSources” parameter with the path to DZI image file, you must configure the OSD viewer “tileSources” as the new derived IrisTileSource class and assign that class the “serverUrl” parameter with the Iris RESTful server domain as well as the IFE encoded slide file name (excluding the ‘.iris’ extension). See example HTML in Code listings 3 for a simple functional example web-viewer application.
@@
// Otherwise you must download a copy and assign that source
// src="openseadragon/openseadragon.min.js"
@@
3.3 Error Handling
The server is runtime hardened with numerous try-catch blocks that prevent runtime errors from escaping each stack. In the event of invalid requests or server-side errors, the server will terminate parsing only a single request and immediately respond with a series of HTTP error codes and response body narratives explaining the server error if such a typical error exists within the dictionary of known issues (such as failed validation or request range violations). The stack thread will then simply move on in the queue. The major response error categories are available within the IrisRestfulTypes.hpp class definitions. All error response descriptions are returned in the HTTP response body in plain-text form for ease of parsing.
@@ 4.1 Materials and Methods
Illustrative performance testing was conducted for both a single-instance local network deployment (Figure 3, right) and the public facing Iris RESTful Example domain in comparison to other well-established and publicly available WSI servers, specifically Path-Presenter and the National Cancer Institute (NCI) imaging data commons (IDC) Slim viewer[6]. The slides were selected for similar imaging characteristics (JPEG encoded 256 x 256 pixel tiles) which resulted in similar packet sizes. Iris Restful Example did not use a content delivery network (CDN); although we are unable to verify that CDN caching was not used by Slim or PathPresenter. Performance was evaluated using Locust[14], a widely accepted open-source and flexible load testing tool. Locust explicitly does not allow local data caching during a test. A ramped load test was chosen to allow for automatic instance scaling in the publicly facing WSI servers. Users were scaled up to 160 and 220 concurrent users for public WSI server and local network tests, respectively. Each virtual user issued a tile request at 1-2 ms intervals after receiving the prior response. Custom locust python scripts were created for each target server and tiles were randomly sampled from within image bounds. These scripts have been made available, along with the raw results of these tests (see Data Availability section). All tests were performed on a 2020 13-in M1 MacBook Pro (Apple, Palo Alto, CA) with 8 GB of RAM running macOS 15.4, with Locust scripts executed using Python 3.11 in an Anaconda environment over a $\geq$ 200 mbps wireless (Wifi) connection.
The single-instance local network deployment was hosted on Turing RTK1 (Turing Machines Inc, Cupertino, CA) 8-core Rockchip RK3588 system on module (SoM) circuit board with 8 GB of RAM running Ubuntu 24.04 within a Docker container and with a 1 gigabit Ethernet switch. The Turing RTK1 SoM is similar to the Raspberry Pi CM4 or the Jetson Orin Nx. The publicly facing Iris RESTful Example domain is hosted on AWS using an application load balancer in front of a Fargate elastic container service with 1 virtual CPU and 2 GB of RAM and set to scale-out at either 70% CPU consumption or greater than 500 requests/sec per instance (up to 100 USD/month with aggressive load testing). Both the local network deployment and Iris RESTful Example domain served IFE formatted tiles using the Iris RESTful API. Path-Presenter served DZI formatted tiles using a static HTTP server without an API. The NCI IDC repository served DICOM formatted frames using the WADO-RS API.
@@ 5 Discussion
We previously published on the rapid encoding speeds of IFE files from source WSI files such as DICOM due to the highly parallelized encoding capabilities afforded by the layer-agnostic tile ordering within the file specification[10]. This aids in the value of the Iris RESTful server by allowing rapid conversion from source DICOM to IFE files, which may then be served to OSD viewers. The current versions of the Iris Codec IFE libraries can generate a derived pyramid file (with 2x or 4x downsampling) from a single highest-resolution DICOM or vendor format source within 1-2 mintues – or much faster if OpenSlide decoding[9] is not needed. We continue to work on improving encoding systems and in future releases will add stream encoding (a slide scanner API) to allow for simultaneous generation of archive DICOM and IFE files directly from the scanner to make IFE slides available to Iris RESTful in WSI workflows immediately. The IFE is in the public domain under a Creative Commons Attribution-No Derivative 4.0 license, and Iris Codec and IrisRESTful server implementations are available to the community under the MIT software license. The IrisTileSource is a part of the OSD repository and follows the OSD ‘s BSD-3 license.
In this paper we performed stressful load testing on WSI systems to assess system durability and performance when these servers. This includes both expected daily loads early in the load test, as well as abnormally high volumes of requests. The early metrics better align with what we have previously observed with the Slim viewer of an approximately 200 ms median response time (Figure 4, left-side) in our prior publication[10]. The Iris RESTful Example domain response times of approximately 20-30 ms also align closely with what we experienced when using our example domain during the development of the IrisTileSource and example browser-based WSVs.
@@ 6 Conclusion
The Iris RESTful Server and OSD IrisTileSource allow for immediate and easy replacement of DZI images with IFE encoded slide files within digital pathology WSI workflows. Prepackaged container images allow for server deployment in a scalable microservice architecture and require altering only four lines of viewer code to begin rendering with the new IrisTileSource within existing OSD viewers. Written in C++ with the highly templated Boost Beast library and on top of the IFE deserialization library this is a low-overhead fast deployment that can issue slide tiles in 21 ms on average under an high request load of over 3883 tile requests per second, even when hosted on a small \$200 USD compute module chipboard (the equivalent of a Raspberry-Pi CM4). We provide the Iris RESTful Example domain, a server that hosts example IFE encoded slide files and responds natively to Iris RESTful API calls, for community use to evaluate HTTPS implementations with the Iris RESTful API and accelerate the development process. All relevant Iris implementations are available under the MIT software license.